# RLeXplore: Accelerating Research in Intrinsically-Motivated Reinforcement Learning

### v1

@@ Abstract
Extrinsic rewards can effectively guide reinforcement learning (RL) agents in specific tasks. However, extrinsic rewards frequently fall short in complex environments due to the significant human effort needed for their design and annotation. This limitation underscores the necessity for intrinsic rewards, which offer auxiliary and dense signals and can enable agents to learn in an unsupervised manner. Although various intrinsic reward formulations have been proposed, their implementation and optimization details are insufficiently explored and lack standardization, thereby hindering research progress. To address this gap, we introduce RLeXplore, a unified, highly modularized, and plug-and-play framework offering reliable implementations of eight state-of-the-art intrinsic reward algorithms. Furthermore, we conduct an in-depth study that identifies critical implementation details and establishes well-justified standard practices in intrinsically-motivated RL. The source code for RLeXplore is available at https://github.com/RLE-Foundation/RLeXplore.
11footnotetext: Equal Contribution.
@@ 1 Introduction
Reinforcement learning (RL) provides a framework for training agents to solve tasks by learning from interactions with an environment. At the core of RL is the optimization of a reward function, where agents aim to maximize cumulative rewards over time [1].
Figure 1: Episode Return achieved by the intrinsic rewards in RLeXplore in SuperMarioBros.
@@
However, in complex environments, defining extrinsic rewards that effectively guide an agent’s learning process can be impractical, often requiring domain-specific expertise. In practice, poorly defined extrinsic rewards can lead to sparse-reward settings, where RL agents struggle due to the lack of a meaningful learning signal [2]. As the RL community tackles increasingly complex problems, such as training generally capable RL agents, there is a need for more autonomous agents capable of learning valuable behaviors without relying on dense supervision [3]. To address this challenge, the concept of intrinsic rewards has emerged as a promising approach in the RL community [4, 5, 6, 7, 8, 9]. Intrinsic rewards provide agents with additional learning signals, enabling them to explore and acquire skills across diverse environments beyond what extrinsic rewards alone can offer.
@@
However, computing intrinsic rewards often requires learning auxiliary models, heavy engineering and performing expensive computations, making reproducibility challenging. While several formulations of intrinsic rewards have been proposed [5, 7, 10], each with its potential benefits for improving agent learning, the field lacks a comprehensive understanding of the comparative advantages and challenges posed by these formulations. Importantly, existing literature reports varying performance when using the same intrinsic rewards, reinforcing the need for a standardized framework and a deeper understanding of the optimization and implementation details.
@@
In this paper, we introduce RLeXplore, an open-source library containing high-quality implementations of state-of-the-art intrinsic rewards. Our work presents a systematic study aimed at addressing gaps in understanding the critical implementation and optimization details of intrinsic rewards. To guide our investigation, we formulate numerous questions, aiming to uncover the intricacies of intrinsic rewards and their impact on RL agent performance. Our results highlight the importance of thoughtful implementation design for intrinsic rewards, showing that naive implementations can lead to suboptimal performance. Through carefully studied design decisions, we demonstrate significant performance gains.
Our contributions are threefold. Firstly, we provide a high-quality open-source repository for training RL agents, featuring the implementation of the most widely recognized intrinsic rewards. Secondly, we present a systematic evaluation that identifies the key implementation and optimization details critical to the success of intrinsic reward methods in RL. Lastly, we provide a comparative analysis of the agents’ performance across challenging environments, establishing a foundation for future research in intrinsically-motivated RL.
@@ 2 Related Work
Intrinsic rewards often rely on heavily engineered implementations to stabilize their optimization. The latter difficulties reproducibility and causes different works in the literature to report varying performance in popular environments. Some works have benchmarked intrinsic rewards in specific environments [11, 12, 10], yet they do not provide details on the importance of the design decisions in the implementation and optimization of the intrinsic rewards. In this work, we introduce RLeXplore, a more comprehensive framework that contains the most widely-used intrinsic rewards and provides the RL community with a unified framework to accelerate research and compare baselines in intrinsically-motivated RL. In the following, we overview existing formulations for intrinsic rewards of different natures and introduce the methods included in RLeXplore.
@@ 2.1 Count-Based Exploration
Count-based exploration methods provide intrinsic rewards by measuring the novelty of states, defined to be inversely proportional to the state visitation counts [13, 14, 15, 16]. In finite state spaces, count-based methods perform near optimally [13]. For this reason, these methods have been established as appealing techniques for driving structured exploration in RL. However, they do not scale well to high-dimensional state spaces [17, 18]. Pseudo-counts provide a framework to generalize count-based methods to high-dimensional and partially observed environments [17, 19, 20]. [4] proposed random network distillation (RND), which uses the prediction error against a fixed network as a learning signal that is correlated to counts. Recently, [8] used an elliptical bonus (E3B) and showed that such an objective provides a generalization of counts to high-dimensional spaces. In RLeXplore, we include Pseudo-counts, RND, and E3B as representatives of the state-of-the-art count-based methods.
@@ 2.2 Curiosity-Driven Exploration
Curiosity-based objectives train agents to interact with the environment seeking to experience outcomes that are not aligned with the agents’ predictions [21]. Hence, curiosity-driven exploration usually involves training an agent to increase its knowledge about the environment (e.g., environment dynamics) [22, 5, 23]. The intrinsic curiosity module (ICM) [5, 2] learns a joint embedding space with inverse and forward dynamics losses and was the first curiosity-based method successfully applied to deep RL settings. Disagreement [9] further extended ICM by using the variance over an ensemble of forward-dynamics models to compute curiosity. However, curiosity-driven methods are consistently found to be unsuccessful when the environment has irreducible noise [24]. To address the problem, [6] proposed RIDE, which uses the difference between two consecutive state embeddings as the intrinsic reward and encourages the agent to choose actions that result in significant state changes. In general, curiosity-based objectives remain amongst the most popular intrinsic rewards in deep RL applications to this day. In RLeXplore, we include ICM, Disagreement, and RIDE as representatives of the state-of-the-art curiosity-driven methods.
@@ 2.3 Global and Episodic Exploration
Towards more general and adaptive agents, recent works have studied decision-making problems in contextual Markov decision processes (MDPs) (e.g., procedurally-generated environments) [6, 8, 25]. Contextual MDPs require episodic-level exploration, where novelty estimates are reset at the beginning of each episode. [26] showed that both global and episodic exploration modalities have unique benefits and proposed combined objectives that achieve remarkable performance across many MDPs of different structures. NGU [7] and RIDE [6] also instantiate both global and episodic bonuses. Inspired by this recent line of works, in this paper, we study novel combinations of objectives for exploration that achieve impressive results in contextual MDPs.
@@ 3 Background
We frame the RL problem considering a MDP [27, 28] defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},E,P,d_{0},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $E:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the extrinsic reward function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition function that defines a probability distribution over $\mathcal{S}$, $d_{0}\in\Delta(\mathcal{S})$ is the distribution of the initial observation $\bm{s}_{0}$, and $\gamma\in[0,1)$ is a discount factor. The goal of RL is to learn a policy $\pi_{\bm{\theta}}(\bm{a}|\bm{s})$ to maximize the expected discounted return:
$J_{\pi}(\bm{\theta})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}E_{t}\right].$ (1)
Intrinsic rewards augment the learning objective to improve exploration. Letting $I:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denote the intrinsic reward function, the augmented optimization objective is:
$J_{\pi}(\bm{\theta})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(E_{t}+\beta_{t}\cdot I_{t})\right],$ (2)
where $\beta_{t}=\beta_{0}(1-\kappa)^{t}$ controls the degree of exploration, and $\kappa$ is a decay rate.
@@ 4 RLeXplore
In this section, we present RLeXplore, a unified, highly-modularized and plug-and-play framework that currently provides high-quality and reliable implementations of eight state-of-the-art intrinsic reward algorithms***RLeXplore complies with the MIT License.. Comparing multiple intrinsic reward methods under fair conditions is challenging due to various confounding factors, such as using distinct backbone RL algorithms (e.g., PPO [29], DQN [30], IMPALA [31]), optimization (e.g., reward and observation normalization, network architecture) and evaluation details (e.g., environment configuration, algorithm hyperparameters). RLeXplore is designed to provide a unified framework with standardized procedures for implementing, computing, and optimizing intrinsic rewards.
Figure 2: The workflow of RLeXplore. (a) RLeXplore monitors the agent-environment interactions and obtains data samples via a .watch() function. (b) For the sampled transitions, RLeXplore calculates the corresponding intrinsic rewards via a .compute() function and performs model updating via a .update() function.
@@ 4.1 Architecture
The core design decision of RLeXplore involves decoupling the intrinsic reward modules from the RL optimization algorithms, which enables our intrinsic reward implementations to be integrated with any desired RL algorithm (or existing library, see Appendix C). Figure 2 illustrates the basic workflow of RLeXplore, which consists of two parts: data collection (i.e., policy rollout) and reward computation.
At each time step, the agent receives observations from the environment and predicts actions. The environment then executes the actions and returns feedback to the agent, which consists of a next observation, a reward, and a terminal signal. During the data collection process, the .watch() function is used to monitor the agent-environment interactions. For instance, E3B [8] updates an estimate of an ellipsoid in an embedding space after observing every state. At the end of the data collection rollouts, .compute() computes the corresponding intrinsic rewards. Note that .compute() is only called once per rollout using batched operations, which makes RLeXplore a highly efficient framework. Additionally, RLeXplore provides several utilities for reward and observation normalization. Finally, the .update() function is called immediately after .compute() to update the reward module if necessary (e.g., train the forward dynamics models in Disagreement [9] or the predictor network in RND [4]). Section C illustrates the usage of the aforementioned functions. All operations are subject to the standard workflow of the Gymnasium API [32].
We provide various examples of using RLeXplore with popular RL frameworks (e.g., stable-baselines3 [33] and CleanRL [34]) in Appendix C.

### v2

@@ Abstract
Extrinsic rewards can effectively guide reinforcement learning (RL) agents in specific tasks. However, extrinsic rewards frequently fall short in complex environments due to the significant human effort needed for their design and annotation. This limitation underscores the necessity for intrinsic rewards, which offer auxiliary and dense signals and can enable agents to learn in an unsupervised manner. Although various intrinsic reward formulations have been proposed, their implementation and optimization details are insufficiently explored and lack standardization, thereby hindering research progress. To address this gap, we introduce RLeXplore, a unified, highly modularized, and plug-and-play framework offering reliable implementations of eight state-of-the-art intrinsic reward methods. Furthermore, we conduct an in-depth study that identifies critical implementation details and establishes well-justified standard practices in intrinsically-motivated RL. Our documentation, examples, and source code are available at https://github.com/RLE-Foundation/RLeXplore.
@@ 1 Introduction
Reinforcement learning (RL) provides a framework for training agents to solve tasks by learning from interactions with an environment. At the core of RL is the optimization of a reward function, where agents aim to maximize cumulative rewards over time (Sutton & Barto, 2018). However, in complex environments, defining extrinsic rewards that effectively guide an agent’s learning process can be impractical, often requiring domain-specific expertise. In practice, poorly defined extrinsic rewards can lead to sparse-reward settings, where RL agents struggle due to the lack of a meaningful learning signal (Burda et al., 2019a).
@@
As the RL community tackles increasingly complex problems, such as training generally capable RL agents, there is a need for more autonomous agents capable of learning valuable behaviors without relying on dense supervision (Jiang et al., 2023). To address this challenge, the concept of intrinsic rewards has emerged as a promising approach in the RL community (Burda et al., 2019b; Pathak et al., 2017; Raileanu & Rocktäschel, 2020; Badia et al., 2020; Henaff et al., 2022; Pathak et al., 2019). Intrinsic rewards provide agents with additional learning signals, enabling them to explore and acquire skills across diverse environments beyond what extrinsic rewards alone can offer. However, computing intrinsic rewards often requires learning auxiliary models, heavy engineering, and performing expensive computations, making reproducibility challenging.
@@
While several formulations of intrinsic rewards have been proposed (Pathak et al., 2017; Badia et al., 2020; Laskin et al., 2021), each with its potential benefits for improving agent learning, the field lacks a comprehensive understanding of the comparative advantages and challenges posed by these methods. Importantly, existing literature reports varying performance when using the same intrinsic rewards, reinforcing the need for a standardized framework and a deeper understanding of the optimization and implementation details.
In this paper, we introduce RLeXplore, an open-source library containing high-quality implementations of state-of-the-art (SOTA) intrinsic rewards. RLeXplore offers a plug-and-play framework for researchers working on intrinsically-motivated RL, enabling them to seamlessly integrate intrinsic rewards into their projects. Specifically, RLeXplore (1) facilitates fair comparisons across multiple intrinsic reward methods, (2) can be easily integrated with various RL frameworks, and (3) streamlines the development of new intrinsic reward methods. In Table 1, we compare the performance of the implementations in RLeXplore with the original results reported in previous works. In Appendix D, we provide the full details on reproducibility with RLeXplore.
Table 1: Summary of comparative results from our RLeXplore implementations and prior work. Right two columns report success rates (of episodes solving the task), except for Procgen where we report mean episode rewards after training. See Appendix D for reproducibility and evaluation details.
Environment Intrinsic Reward Original RLeXplore
SuperMarioBros (10M Steps) RIDE $23\%$ $\bm{50\%}$
SuperMarioBros (10M Steps) ICM $30\%$ $30\%$
MiniGrid-DoorKey-16×16 (10M Steps) ICM $0\%$ $\bm{60\%}$
MiniGrid-DoorKey-16×16 (10M Steps) RND $0\%$ $\bm{60\%}$
[…]
MiniGrid-DoorKey-8×8 (1M Steps) RE3 $50\%$ $\bm{95\%}$
MiniGrid-DoorKey-8×8 (1M Steps) RND $0\%$ $\bm{82\%}$
MiniGrid-DoorKey-8×8 (1M Steps) ICM $20\%$ $\bm{83\%}$
Procgen - 200 Mazes (25M Steps) E3B $3.00$ $\bm{4.10}$
Procgen - 200 Mazes (25M Steps) ICM $2.50$ $\bm{5.90}$
Procgen - 200 Mazes (25M Steps) RND $1.70$ $\bm{5.00}$
To support these capabilities, we have provided extensive documentation that includes detailed guides on using RLeXplore, along with comprehensive code tutorials. These resources are designed to make it straightforward for users to get started with RLeXplore, regardless of their prior experience with intrinsic rewards in RL. In Appendix D.6, we provide an overview of the main differences and advantages of RLeXplore compared to existing RL libraries.
We aim for the community to adopt RLeXplore as a standard tool for evaluating intrinsic reward methods, reducing implementation efforts, and mitigating inconsistencies in results and conclusions.
@@
Our work presents a systematic study aimed at addressing gaps in understanding the critical implementation and optimization details of intrinsic rewards. We investigate the design of different intrinsic reward methods and (1) highlight challenges in the reproducibility of prior work, and (2) share highly performant reimplementations of many popular methods. To guide our investigation, we formulate numerous questions, aiming to uncover the intricacies of intrinsic rewards and their impact on RL agent performance. Our results highlight the importance of thoughtful implementation design for intrinsic rewards, showing that naive implementations can lead to suboptimal performance. Through carefully studied design decisions, we demonstrate significant performance gains.
Our results show that with RLeXplore, RL agents can learn emergent behaviours autonomously, solving multiple levels of SuperMarioBros without task rewards. Additionally, we show that intrinsic rewards enable RL agents to obtain great performance on complex sparse-reward tasks like Procgen-Maze, MiniGrid, the ALE-5 hard-exploration tasks and Ant-UMaze.
@@ 2 Related Work
While some works have benchmarked intrinsic rewards in specific environments (Taiga et al., 2021; Wang et al., 2022; Laskin et al., 2021), their lack of detailed discussions on implementation and optimization leads to reproducibility problems (Voelcker et al., 2024). In this work, we introduce RLeXplore, a comprehensive framework that incorporates the most widely-used intrinsic rewards, which provides a standardized approach to enhance reproducibility, accelerate research, and facilitate the comparison of baselines in intrinsically-motivated RL. In the following, we overview existing formulations for intrinsic rewards of different natures and introduce the methods included in RLeXplore.
@@ 2.1 Count-Based Exploration
Count-based exploration methods provide intrinsic rewards by measuring the novelty of states, defined to be inversely proportional to the state visitation counts (Strehl & Littman, 2008; Tang et al., 2017; Machado et al., 2020; Jo et al., 2022). In finite state spaces, count-based methods perform near optimally (Strehl & Littman, 2008). For this reason, these methods have been established as appealing techniques for driving structured exploration in RL. However, they do not scale well to high-dimensional state spaces (Bellemare et al., 2016; Lobel et al., 2023). Pseudo-counts provide a framework to generalize count-based methods to high-dimensional and partially observed environments (Bellemare et al., 2016; Ostrovski et al., 2017; Martin et al., 2017). Burda et al. (2019b) proposed random network distillation (RND), which uses the prediction error against a fixed network as a learning signal that is correlated to counts. Recently, Henaff et al. (2022) proposed E3B and showed that the intrinsic objective provides a generalization of counts to high-dimensional spaces. In RLeXplore, we include Pseudo-counts, RND, and E3B as representatives of the state-of-the-art count-based methods.
@@ 2.2 Curiosity-Driven Exploration
Curiosity-based objectives train agents to interact with the environment seeking to experience outcomes that are not aligned with the agents’ predictions (Aubret et al., 2023). Hence, curiosity-driven exploration usually involves training an agent to increase its knowledge about the environment (e.g., environment dynamics) (Stadie et al., 2015; Pathak et al., 2017; Yu et al., 2020). The intrinsic curiosity module (ICM) (Pathak et al., 2017; Burda et al., 2019a) learns a joint embedding space with inverse and forward dynamics losses and was the first curiosity-based method successfully applied to deep RL settings. Disagreement (Pathak et al., 2019) further extended ICM by using the variance over an ensemble of forward-dynamics models to compute curiosity. However, curiosity-driven methods are consistently found to be unsuccessful when the environment has irreducible noise (Savinov et al., 2019). To address the problem, Raileanu & Rocktäschel (2020) proposed RIDE, which uses the difference between two consecutive state embeddings as the intrinsic reward and encourages the agent to choose actions that result in significant state changes. In general, curiosity-based objectives remain amongst the most popular intrinsic rewards in deep RL applications to this day. In RLeXplore, we include ICM, Disagreement, and RIDE as representatives of the state-of-the-art curiosity-driven methods.
@@ 2.3 Global and Episodic Exploration
Towards more general and adaptive agents, recent works have studied decision-making problems in contextual Markov decision processes (MDPs) (e.g., procedurally-generated environments) (Raileanu & Rocktäschel, 2020; Henaff et al., 2022; Matthews et al., 2024). Contextual MDPs require episodic-level exploration, where novelty estimates are reset at the beginning of each episode. Henaff et al. (2023) showed that both global and episodic exploration modalities have unique benefits and proposed combined objectives that achieve remarkable performance across many MDPs of different structures. NGU (Badia et al., 2020) and RIDE (Raileanu & Rocktäschel, 2020) also instantiate both global and episodic bonuses. Inspired by this recent line of work, in this paper, we study novel combinations of objectives for exploration that achieve impressive results in contextual MDPs.
@@ 3 Background
We frame the RL problem considering a MDP (Bellman, 1957; Kaelbling et al., 1998) defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},R,P,d_{0},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition function that defines a probability distribution over $\mathcal{S}$, $d_{0}\in\Delta(\mathcal{S})$ is the distribution of the initial observation $\bm{s}_{0}$, and $\gamma\in[0,1)$ is a discount factor. The goal of RL is to learn a policy $\pi_{\bm{\theta}}(\bm{a}|\bm{s})$ to maximize the expected discounted return:
$J_{\pi}(\bm{\theta})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t}\right].$ (1)
Following prior work in intrinsically-motivated RL (Burda et al., 2019b), we consider an augmented reward function that combines both extrinsic and intrinsic components. The overall reward signal at time step $t$ is defined as:
$R^{\rm total}_{t}=R_{t}+\beta_{t}\cdot I_{t},$ (2)
where $I:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the intrinsic reward, $\beta_{t}=\beta_{0}(1-\kappa)^{t}$ is the exploration coefficient controlling the relative weight of intrinsic motivation over time, and $\kappa$ is a decay rate. This formulation enables the agent to balance task-specific objectives with exploratory behavior, particularly in sparse or deceptive reward settings. Finally, the augmented optimization objective is:
$J_{\pi}(\bm{\theta})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(R_{t}+\beta_{t}\cdot I_{t})\right].$ (3)
In Appendix A, we present a detailed overview of the SOTA intrinsic reward methods that we implement in RLeXplore.
(a) Conceptual framework.
(b) Implementation framework.
(c) Create mixed intrinsic rewards using the Fabric class.
Figure 1: The workflow of RLeXplore. (a) RLeXplore provides a decoupled module for intrinsic rewards that integrates seamlessly with the RL training loop. RLeXplore implements 8 SOTA intrinsic rewards and adapts to the unmodified RL training loop. (b) RLeXplore monitors the agent-environment interactions and gathers data samples using the .watch() function. After collecting experience rollouts, RLeXplore computes the corresponding intrinsic rewards using the .compute() function and updates the auxiliary models via the .update() function. (c) RLeXplore provides a Fabric class that allows developers to combine multiple intrinsic rewards in an elegant manner. In Appendix C.6 we provide more details on how to add new intrinsic rewards to RLeXplore.
@@ 4 RLeXplore
In this section, we present RLeXplore, a unified, highly-modularized and plug-and-play framework that currently provides high-quality and reliable implementations of eight SOTA intrinsic reward methods [^1] RLeXplore complies with the MIT License.. Comparing multiple intrinsic reward methods under fair conditions is challenging due to various confounding factors, such as using distinct RL frameworks (e.g., PPO (Schulman et al., 2017), DQN (Mnih et al., 2013), IMPALA (Espeholt et al., 2018)), optimization (e.g., reward and observation normalization, network architecture) and evaluation details (e.g., environment configuration, algorithm hyperparameters). RLeXplore is designed to provide a unified framework with standardized procedures for implementing, computing, and optimizing intrinsic rewards.
@@ 4.1 Architecture
The core design decision of RLeXplore involves decoupling the intrinsic reward modules from the RL optimization algorithms, which enables our intrinsic reward implementations to be integrated with any desired RL algorithm (or existing library, see Appendix C and the official integration examples). Figure 1 illustrates the basic workflow of RLeXplore, which consists of two parts: data collection (i.e., policy rollout) and reward computation.
Commonly, at each time step, the agent receives observations from the environment and predicts actions. The environment then executes the actions and returns feedback to the agent, which consists of a next observation, a reward, and a terminal signal. During the data collection process, the .watch() function is used to monitor the agent-environment interactions. For instance, E3B (Henaff et al., 2022) updates an estimate of an ellipsoid in an embedding space after observing every state. At the end of the data collection rollouts, .compute() computes the corresponding intrinsic rewards. Note that .compute() is only called once per rollout using batched operations, which makes RLeXplore a highly efficient framework. Additionally, RLeXplore provides several utilities for reward and observation normalization. Finally, the .update() function is called immediately after .compute() to train the reward module if necessary (e.g., train the forward dynamics models in Disagreement (Pathak et al., 2019) or the predictor network in RND (Burda et al., 2019b)). Appendix C illustrates the usage of the aforementioned functions. All operations are subject to the standard workflow of the Gymnasium API (Towers et al., 2023).
In particular, recent research (Henaff et al., 2023) has highlighted that mixed intrinsic rewards can significantly promote the agent’s exploration capability by providing comprehensive exploration incentives. In RLeXplore, we provide a Fabric class that allows developers to combine multiple intrinsic rewards in an elegant manner, as illustrated in Appendix C.3.
RLeXplore offers several benefits to the research community:
  • For researchers seeking reliable tools for benchmarking and general applications: RLeXplore provides high-quality implementations of popular intrinsic reward methods, useful in both research and practical applications. It can be seamlessly integrated with existing RL libraries. We provide specific examples of integrating RLeXplore with Stable-Baselines3 (Raffin et al., 2021), CleanRL (Huang et al., 2022b), and RLLTE (Yuan et al., 2023) in Appendix C, F, and G.
  • For developers experimenting with new intrinsic rewards: RLeXplore offers modular components, such as various embedding networks and a standardized workflow. This setup facilitates the creation, modification, and testing of new ideas. Detailed examples are available in the code repository and documentation.
  • For promoting collaboration and accelerating progress: We have published a space using Weights & Biases (W&B) to store reusable experiment results on recognized benchmarks. This initiative aims to enhance collaboration within the research community and speed up progress by providing easy access to established benchmark results.